OpenStudy (anonymous):

Given the arithmetic sequence an = 6 − 4(n + 2), what is the domain for n? @kc_kennylau

OpenStudy (kc_kennylau):

\[n\in\mathbb N, n=[1,+\infty)\]

OpenStudy (anonymous):

huh lol english pls

OpenStudy (kc_kennylau):

for all natural numbers excluding 0 :D

OpenStudy (anonymous):

so all real numbers?

OpenStudy (kc_kennylau):

no, just positive integers

OpenStudy (kc_kennylau):

\[n\in\mathbb Z^+\]

OpenStudy (anonymous):

Thank You!

OpenStudy (kc_kennylau):

no problem :)

OpenStudy (agent0smith):

Basically \[\large n \ge 1\]